OpenStudy (anonymous):

4(x-1)=x+11 help!!! I have homework

OpenStudy (plasmataco):

how in summer, that SUCKS

OpenStudy (anonymous):

School probably started

OpenStudy (plasmataco):

First get rid of the parentheses by multiplying 4 with x-1

OpenStudy (plasmataco):

U get 4x-4

OpenStudy (plasmataco):

Now move the x's to one side and numbers to the other side

OpenStudy (plasmataco):

4x-4=x+11 turns into 4x-x=11+4

OpenStudy (plasmataco):

U get 3x=15... So x=5

OpenStudy (anonymous):

Yea in high school doin algebra schools just started anyway my friends need to help me so let's get down to business the question is 4(x-1)=x + 11

OpenStudy (plasmataco):

I just solveed it... Look above. :3

OpenStudy (rushwr):

Hey this easy ! \[4(x-1) = x+11\] \[4x-4=x+11\] \[4x-x=11+4\] \[3x=15\] \[x=\frac{ 15 }{ 3}\] \[x=5\]

OpenStudy (anonymous):

Rush made it look easy lol

OpenStudy (plasmataco):

Lol... He did the exact same thing as I did :3

OpenStudy (anonymous):

Lol but she showed her Work literally

OpenStudy (plasmataco):

That's not fair... I'm working on an iPad.

OpenStudy (anonymous):

Not saying u didn't do a good job

OpenStudy (plasmataco):

I can't upload pictures, which rly stinks

OpenStudy (anonymous):

Yea it does

OpenStudy (anonymous):

Hay step 3 is -5m/_5 but It equals m=-3 make sure you can divide simplify

OpenStudy (rushwr):

guys chill ! Hold on @Plasmataco

OpenStudy (anonymous):

Haha sirry

OpenStudy (anonymous):

Sorry

OpenStudy (plasmataco):

Lol sry.

OpenStudy (plasmataco):

No need to give me a medal, it's fine

OpenStudy (rushwr):

no nw it's done so PEACE !

OpenStudy (anonymous):

U guys u were rite but u didn't simplify

OpenStudy (plasmataco):

@prettyprincess1 asked u a question I think

OpenStudy (plasmataco):

@Rushwr

OpenStudy (anonymous):

U oh way @plasmataco

OpenStudy (plasmataco):

?

OpenStudy (rushwr):

Actually the answer is x=5

OpenStudy (plasmataco):

Don't ask me, I'm only 13

OpenStudy (anonymous):

Thanks beasties I love y'all the best ok fan me!

OpenStudy (anonymous):

Thank u